RC Meeting, Thursday, October 28 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. Beam delivery back to Halls B,C,D on owl shift.

Hall C Status

Continue production at 42 muA.

  1. day shift: productive, 160 good pions recorded.
  1. swing shift: Empty target run #12863. A total (runs 12860-12864) of 111 good pions, including a grand total of one on the dummy target run.
  1. owl shift: very productive. 30 min lost to Acc. QE measurements.

Number of good pions is 5764+.

Current Run Plan

  1. BCM calibration: Thursday, @ 2:30 pm. As per the 8:00 am meeting this morning we should be OK to get up to the desired 60 muA so we can keep the errors on the fitted gains small.
  2. Continue data taking until early Nov

Current Issues

  1. Given the uncertainty associated with recent Hall A events it is not clear that the maintenance and retuning will happen as scheduled (Tu and Th, respectively). Need to follow this up with the scheduling committee.
  2. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  3. HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
  4. Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
  5. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  6. Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
  7. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links